OpenStudy (anonymous):

Jimmy needs to pass a college class. On his first exam he got a 44% and on his 2nd exam he didn't do any better and got a 66%. What percent does jimmy need to get on his last exam so that he can pass with a C.

OpenStudy (anonymous):

44/100x-->first exam 66/100x-->second exam

OpenStudy (anonymous):

what am I doing with this?

OpenStudy (anonymous):

solving for x?

OpenStudy (anonymous):

Well i cant really solve for x now can i

OpenStudy (anonymous):

Yea I don't think that is it

OpenStudy (anonymous):

yes u are right

OpenStudy (anonymous):

and how are you getting 110?

OpenStudy (anonymous):

oh alright so then

OpenStudy (anonymous):

i think 88/100

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

I feel as if there is something im not seeing

OpenStudy (anonymous):

was 88/100 not the answer?

OpenStudy (anonymous):

i was not confident with my answer.i think other way also 44%-x=56% and 66%-x=34% so total x=90%.i was not sure .whether my answer is correct

OpenStudy (anonymous):

first 44/100x so 11/25

OpenStudy (anonymous):

well i do know that 44%-x=56% equals -3/25 while 66%-x=34% equal 8/25

OpenStudy (anonymous):

second 33/50

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

i think this will be the right answer 90%

OpenStudy (anonymous):

see in first exam 11/25 so left is 14/25 and in second 33/50 and left is 17/50 and addin both we get 45/50*100=90

OpenStudy (anonymous):

14/25+17/50=45/50 so 45/50x*100=90

OpenStudy (anonymous):

is it correct

OpenStudy (anonymous):

probably a 77%

OpenStudy (anonymous):

@dumbcow

OpenStudy (anonymous):

@ganeshie8 answer for oldschooldexter

OpenStudy (dumbcow):

assuming equal weights for the exams and 70 as passing with a C --> (44+66+X)/3 = 70 44+66+X = 210 X = 100

OpenStudy (anonymous):

so c=100%

OpenStudy (dumbcow):

yeah

OpenStudy (anonymous):

if we take 100 then x=210

OpenStudy (anonymous):

(44+66+x)/3=100 means x=210

OpenStudy (dumbcow):

hmm no x=190 in that case 66+44 = 110 110 + 190 = 300

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x=190 also correct?

OpenStudy (anonymous):

bit confusing from yesterday.i want to put him correct solution

OpenStudy (dumbcow):

well like i said assuming the class grade consists of only the 3 equally weighted exams then its simply the average of the 3 exam scores --> (44+66+X)/3 = class grade if he wants at least a 70% or C then X = 100%, he needs a perfect score on final exam

OpenStudy (anonymous):

i dont know the correct solution.but it makes a sense

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

help him